OpenStudy (anonymous):

I have no idea were to begin If the graph of an equation of the form y=ab^x goes through (2,18)(5,60.75) then a= b= Help!!!!!

OpenStudy (anonymous):

you'll need to plug both those points into the equation to get two equations with two variables.... then solve the system...

OpenStudy (anonymous):

*solve for a, b

OpenStudy (anonymous):

Let's plug that first point into the equation: (2, 18) into y=ab^x becomes \(\large 18=a\cdot b^2 \) this is one of your equations... now do the same for the other point...

OpenStudy (anonymous):

5=a*b^60.75 is this right

OpenStudy (anonymous):

i think you mixed up your x and y there...

OpenStudy (anonymous):

60.75=a*b^5 now good

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

now you have two equations with two unknowns.... you can solve this system by substitution....

OpenStudy (anonymous):

Ok by doin?

OpenStudy (anonymous):

substitution.... try solving for a in the first equation and plug that into the second....

OpenStudy (anonymous):

in the first equation, \(\large 18=ab^2 \rightarrow a=\frac{18}{b^2} \) plug this expression for a into the second equation....

OpenStudy (anonymous):

ok I got what you just posted when I did my math. What do I do with the b^5 when plugging into the second equation when solving

OpenStudy (anonymous):

you need to simplify the equation you now have.... what is the equation you have?

OpenStudy (anonymous):

60.75= (18/b^2)*b^5

OpenStudy (anonymous):

ok... simplify that...

OpenStudy (anonymous):

42.75=b^7

OpenStudy (anonymous):

\(\large 60.75=\frac{18}{\cancel{b^2}^1}\cdot \cancel{b^5}b^3 \) \(\large 60.75=18\cdot b^3 \)

OpenStudy (anonymous):

oh ok so then now it is 42.75=b^3

OpenStudy (anonymous):

how did you get 42.75?

OpenStudy (anonymous):

60.75-18=42.75

OpenStudy (anonymous):

no... you should divide...

OpenStudy (anonymous):

3.375=b^3 now

OpenStudy (anonymous):

good... do you know what to do from here?

OpenStudy (anonymous):

3.375^(1/3)=1.5 b=1.5

OpenStudy (anonymous):

good.... now find a...

OpenStudy (anonymous):

remember, a = 18/(b^2)

OpenStudy (anonymous):

so a=8

OpenStudy (anonymous):

good.... now can you write the exponential equation for me?

OpenStudy (anonymous):

y=(8)1.5^x

OpenStudy (anonymous):

nicely done!!! good work.... :)

OpenStudy (anonymous):

thank you and thak you for all your help:)

OpenStudy (anonymous):

yw... and one more thing... if you are uncertain if your answer is correct or not, plug in those points again into the equation you now have... i think you'll see that the answer is correct... bye...